# Limited Memory LRSGA Optimizer to competitive optimization

### v1

@@ Abstract
We introduce LM-LRSGA, a limited-memory variant of Low-Rank Symplectic Gradient Adjustment (LRSGA) for differentiable games. It is an iterative scheme for approximating Nash equilibria at first-order cost while preserving the stabilizing benefits of second-order information. By storing only a limited history of curvature pairs, LM-LRSGA is well suited to high-parameter competitive models such as GANs.
Keywords Nash equilibria $\cdot$ Competitive optimization $\cdot$ GreenAI $\cdot$ Limited memory $\cdot$ Spectral analysis
@@ 1 Introduction
$H(w)=\begin{pmatrix}\partial^{2}_{xx}f(w)&\partial^{2}_{xy}f(w)\\ \partial^{2}_{yx}g(w)&\partial^{2}_{yy}g(w)\end{pmatrix}$,
@@
We address this scalability bottleneck with Limited-Memory Low-Rank Symplectic Gradient Adjustment (LM-LRSGA). The method combines symplectic gradient corrections with L-BFGS-style recursions: instead of storing dense factors, it keeps a short history of curvature pairs capturing essential information from recent iterates. Modified two-loop recursion algorithms compute the LRSGA products implicitly, avoiding matrix formation or storage. Thus LM-LRSGA preserves LRSGA’s stabilizing effect while using linear memory and first-order-like per-step complexity, making it suitable for large-scale adversarial/stochastic settings (e.g., GANs with millions of parameters).
We validate the approach on GAN training over MNIST and Fashion-MNIST, comparing LM-LRSGA against the Adam Kingma and Ba (2014) optimizer. Empirically, LM-LRSGA yields more stable training dynamics and faster convergence, with smooth and synchronized loss trajectories for both generator and discriminator. Quantitatively, it improves Fréchet Inception Distance (FID) by up to $30\%$ over Adam while maintaining comparable computational cost. Sensitivity analyses show robustness to learning rate, history length, and damping factors, highlighting suitability for large-scale and stochastic regimes.
Beyond computation, the method supports sustainable ML: by avoiding explicit second-order computation and reducing memory footprint, LM-LRSGA lowers training time, suggesting efficiency and environmental friendliness.
@@
Section 2 reviews competitive optimization and recalls SGA and LRSGA. Section 3 presents LM-LRSGA, detailing its two-loop recursion algorithms and update rules. Section 4 reports experiments on GANs trained on MNIST and Fashion-MNIST, and presents sensitivity analyses. Section 6 concludes and outlines potential future extensions to large-scale stochastic and distributed settings.
@@ 2 Preliminaries and Mathematical Framework
Solutions to problem 2 are called Nash equilibria (NE) and denoted by $(x^{*},y^{*})$. For a shorter notation, we write $w=(x,y)\in X\times Y$. The first-order stationarity and the second-order necessary conditions for a local NE are:
@@
Moreover, to ensure that the NE is isolated and locally stable, we assume H(w)H(w*) to be invertible and semi-definite positive i.e. $\langle z,\,H(w^{\text{NE}})\,z\rangle\geq 0\quad\text{for all }z\in\mathbb{R}^{m+n}$.
We remark that the game Hessian matrix admits the symmetric/antisymmetric decomposition
$S(x,y)=\frac{H(x,y)+H(x,y)^{\top}}{2},\quad A(x,y)=\frac{H(x,y)-H(x,y)^{\top}}{2}.$
@@
where $I$ represents the identity matrix in the appropriate Euclidean space, $w_{k}:=(x_{k},y_{k})$ the iterate $k$, $\eta>0$ the stepsize and $\tau>0$ controls the strength of the antisymmetric correction.The antisymmetric component of the game Hessian at iterate $w_{k}$ $A(w_{k})$ has the block form:
$A(w_{k})=\begin{pmatrix}0_{m}&\frac{1}{2}\,(\partial_{xy}^{2}f(w_{k})-\partial_{yx}^{2}g(w_{k})^{\top})\\ \frac{1}{2}\,(\partial_{yx}^{2}g(w_{k})-\partial_{xy}^{2}f(w_{k})^{\top})&0_{n}\end{pmatrix}$ (4)
@@
As outlined in the introduction, our recent variant—Low-Rank Symplectic Gradient Adjustment (LRSGA)—proposed in Vater et al. (2025), addresses the high computational cost of explicitly forming the mixed second derivatives $\partial^{2}_{xy}f(x,y)$ and $\partial^{2}_{yx}g(x,y)$ by employing Broyden’s least–change (rank–one) secant updates Broyden (1965) to obtain efficient approximations of these mixed blocks. In particular, we maintain Broyden-type Jacobian estimates with $\mu_{k}\approx\bigl(\partial_{xx}^{2}f\,\big|\,\partial_{xy}^{2}f\bigr)\in\mathbb{R}^{m\times(m+n)}$ and $\nu_{k}\approx\bigl(\partial_{yx}^{2}g\,\big|\,\partial_{yy}^{2}g\bigr)\in\mathbb{R}^{n\times(m+n)}$. Henceforth we retain only the mixed–derivative block–columns, denoted $\mu_{k}]_{2}$ and $\nu_{k}]_{1}$, which approximate $\partial_{xy}^{2}f$ and $\partial_{yx}^{2}g$, respectively.
@@
It is algebraically and computationally equivalent (at the level of per–iteration arithmetic) to (i) form $\mu_{k}$ and $\nu_{k}$ and then extract only the mixed blocks $\mu_{k}]_{2}$ and $\nu_{k}]_{1}$, or (ii) update directly the mixed blocks. Indeed, starting from (5) and partitioning $\mu_{k}=\bigl([\mu_{k}]_{1}\,\big|\,[\mu_{k}]_{2}\bigr)$ and $\nu_{k}=\bigl([\nu_{k}]_{1}\,\big|\,[\nu_{k}]_{2}\bigr)$, one readily verifies
@@
where $f_{k}:=f(w_{k})$ and $g_{k}:=g(w_{k})$. Note that
$(w_{k+1}-w_{k})^{\top}(w_{k+1}-w_{k})=\|x_{k+1}-x_{k}\|_{2}^{2}+\|y_{k+1}-y_{k}\|_{2}^{2}.$
Thus, in either formulation the same quantities—namely the products $\mu_{k}]_{1}(x_{k+1}-x_{k})$ and $\nu_{k}]_{2}(y_{k+1}-y_{k})$—are required, and the per–step computational cost is comparable. Moreover, near a stable Nash equilibria $w^{*}$, the own–player Hessian blocks are assumed to be positive semidefinite by the second–order necessary conditions of Nash equilibria point:
@@
We can assume moreover that these blocks are uniformly bounded in a convex neighborhood $\Omega$ of $w^{*}$:
$0\preceq\partial^{2}_{xx}f(w)\preceq\bar{\lambda}_{x}I_{m},\qquad 0\preceq\partial^{2}_{yy}g(w)\preceq\bar{\lambda}_{y}I_{n},\qquad\forall\,w\in\Omega.$ (7)
(These bounds follow, e.g., from continuity of the Hessians plus compactness of $\Omega$.) Under (7) it is consistent to replace the unknown diagonal blocks by $\varepsilon_{x}I_{m}$ and $\varepsilon_{y}I_{n}$, with
$0\leq\varepsilon_{x}\leq\bar{\lambda}_{x},\qquad 0\leq\varepsilon_{y}\leq\bar{\lambda}_{y},$
which allow us to consider:
$M_{k+1}$ $=M_{k}+\frac{(\partial_{x}f_{k+1}-\partial_{x}f_{k}-\varepsilon_{x}(x_{k+1}-x_{k})-M_{k}(y_{k+1}-y_{k}))(y_{k+1}-y_{k})^{\top}}{(w_{k+1}-w_{k})^{\top}(w_{k+1}-w_{k})}$ (8)
$N_{k+1}$ $=N_{k}+\frac{(\partial_{y}g_{k+1}-\partial_{y}g_{k}-\varepsilon_{y}(y_{k+1}-y_{k})-N_{k}(x_{k+1}-x_{k}))(x_{k+1}-x_{k})^{\top}}{(w_{k+1}-w_{k})^{\top}(w_{k+1}-w_{k})}$
The induced modeling error in the secant residuals remains bounded and can be made small by choosing $\varepsilon_{x},\varepsilon_{y}$ within the spectral bounds in (7). Practically, $\varepsilon_{x},\varepsilon_{y}$ play the role of Tikhonov/Levenberg–Marquardt regularization; setting them to 0 corresponds to ignoring the diagonal blocks, whereas small positive values improve stability without increasing memory.
Therefore, the (slightly revised) LRSGA we employ adopts the following iterative rule:
$w_{k+1}=w_{k}-\eta\,(I-\tau\,\alpha_{k})\,F(w_{k}),$ (9)
where
$\alpha_{k}=\begin{pmatrix}0&\frac{1}{2}(M_{k}-N_{k}^{\top})\\[6.0pt] \frac{1}{2}(N_{k}-M_{k}^{\top})&0\end{pmatrix}\approx A(w_{k}).$ (10)
This work introduces a Limited-Memory LRSGA (LM–LRSGA) variant that preserves the theoretical framework of LRSGA while avoiding the explicit storage of the matrices $M_{k}$ and $N_{k}$. The implementation follows a two-loop recursion scheme in the spirit of L-BFGS Nocedal (1980); Liu and Nocedal (1989), leading to substantial memory savings and making the method suitable for large-scale competitive settings. Further details and evidence are provided in the subsequent sections.
3 Limited Memory LRSGA
@@
As highlighted, LRSGA requires storing the matrices $M_{k}$ and $N_{k}$ to approximate the mixed second derivatives, which can be memory–intensive for high–dimensional models. To address this, we propose a Limited-Memory LRSGA variant that preserves LRSGA’s theoretical robustness while substantially reducing memory usage.
Our strategy follows the two-loop recursion method, successfully employed by L-BFGS Nocedal (1980); Liu and Nocedal (1989). The key idea is to avoid storing the full updated matrices $M_{k}$ and $N_{k}$; instead, we retain only the most recent $\ell$ pairs among the four vectors $\{s_{k}^{x},s_{k}^{y},y_{k}^{f},y_{k}^{g}\}$, defined as
$s_{k}^{x}=x_{k}-x_{k-1},\qquad s_{k}^{y}=y_{k}-y_{k-1},$
$y_{k}^{f}=\partial_{x}f(w_{k})-\partial_{x}f(w_{k-1})-\varepsilon_{x}\,s_{k}^{x},\qquad y_{k}^{g}=\partial_{y}g(w_{k})-\partial_{y}g(w_{k-1})-\varepsilon_{y}\,s_{k}^{y}.$

### v2

# Limited-Memory LRSGA: An Iterative Method for Computing Nash Equilibria in Competitive Optimization Problems

@@ Abstract
We introduce LM-LRSGA, a limited-memory variant of Low-Rank Symplectic Gradient Adjustment (LRSGA) for differentiable games. It is an iterative scheme for approximating Nash equilibria with first-order-like cost while retaining the stabilizing effect of symplectic second-order corrections via low-rank information. By storing only a limited history of curvature pairs, LM-LRSGA is well suited to high-parameter competitive models such as GANs. In particular, we provide a per-iteration spectral stability condition for LRSGA near Nash equilibria, a limited‑memory implementation (LM‑LRSGA) based on adapted two‑loop recursions together with a local convergence analysis for fixed history length, and an empirical evaluation on GAN training on MNIST and Fashion-MNIST, including spectral diagnostics of the training dynamics.
Keywords Nash equilibria $\cdot$ Competitive optimization $\cdot$ GreenAI $\cdot$ Limited memory $\cdot$ Spectral $\cdot$ GANs
@@ 1 Introduction
$H(w)=\begin{pmatrix}D(\partial_{x}f(w))\\ D(\partial_{y}g(w))\end{pmatrix}=\begin{pmatrix}\partial^{2}_{xx}f(w)&\partial^{2}_{xy}f(w)\\ \partial^{2}_{yx}g(w)&\partial^{2}_{yy}g(w)\end{pmatrix}$,
@@
This paper extends LRSGA by adding a spectral stability result, an $\varepsilon$-regularized mixed-block formulation, and a limited-memory variant, Limited-Memory Low-Rank Symplectic Gradient Adjustment (LM-LRSGA), with local convergence analysis. The method combines symplectic gradient corrections with L-BFGS-style recursions: instead of storing dense factors, it keeps a short history of curvature pairs capturing essential information from recent iterates. Modified two-loop recursion algorithms compute the LRSGA products implicitly, avoiding matrix formation or storage. Thus LM-LRSGA preserves LRSGA’s stabilizing effect while using linear memory and first-order-like per-step complexity, making it suitable for large-scale adversarial/stochastic settings (e.g., GANs with millions of parameters).
Unlike L-BFGS, which aims to approximate the full Hessian structure, LM-LRSGA targets only the antisymmetric component responsible for rotational dynamics. As a result, the method avoids the estimation of the full Hessian information, thereby reducing the amount of second-order approximation required.
We validate the approach on GAN training over MNIST and Fashion-MNIST, comparing LM-LRSGA against the Adam Kingma and Ba (2014) optimizer. Empirically, LM-LRSGA yields more stable training dynamics and faster convergence, with smooth and synchronized loss trajectories for both generator and discriminator. Quantitatively, it achieves lower Fréchet Inception Distance (FID) scores than Adam while maintaining comparable computational cost. Sensitivity analyses show robustness to learning rate, history length, and damping factors, highlighting suitability for large-scale and stochastic regimes.
Beyond computation, our method supports sustainable ML by avoiding explicit second-order computation and reducing memory footprint.
@@
Section 2 reviews competitive optimization and recalls SGA and LRSGA.Section 3 establishes a spectral stability condition for LRSGA near Nash equilibria. Section 4 presents LM-LRSGA, detailing its two-loop recursion algorithms and update rules. Section 5 proves local linear convergence of LM-LRSGA under fixed history length and sufficiently small step sizes.Section 6 reports experiments on GANs trained on MNIST and Fashion-MNIST, and presents sensitivity analyses. Section 7 complements the empirical results with spectral and stability diagnostics (Jacobian spectrum estimation and Welch-PSD refinement). Section 8 concludes and outlines future work, including extending LM-LRSGA to larger-scale stochastic and distributed training, to noncooperative games with more than two players, and quantifying the associated computational and carbon footprint.
@@ 2 Preliminaries and Mathematical Framework
Solutions to problem 2 are called Nash equilibria (NE) and denoted by $(x^{*},y^{*})$. For a shorter notation, we write $w=(x,y)\in X\times Y$.
The first-order stationarity and the second-order necessary conditions for a local NE $w^{*}=(x^{*},y^{*})$ are
@@
That is, the game gradient vanishes at $w^{*}$ and each player’s second order derivative with respect to its own variable is positive semidefinite.
We remark that the game Hessian matrix admits the symmetric/antisymmetric decomposition $H(x,y)=S(x,y)+A(x,y)$:
$S(x,y)=\frac{H(x,y)+H(x,y)^{\top}}{2}\quad A(x,y)=\frac{H(x,y)-H(x,y)^{\top}}{2},$
and, in order to guarantee local stability in our analysis, we adopt the sufficient assumption that the symmetric part satisfies $S(w^{*})\succeq\mu I$ for some $\mu>0$. This condition provides a uniform coercive (dissipative) component in the linearized game dynamics, which underpins discrete-time contraction bounds for the preconditioned iteration. In practice, such coercivity is consistent with (and may be reinforced by) common quadratic regularization terms.
@@
where $I$ represents the identity matrix in the appropriate Euclidean space, $w_{k}:=(x_{k},y_{k})$ the iterate $k$, $\eta>0$ the stepsize, $\tau>0$ controls the strength of the antisymmetric correction and $A(w_{k})$ has the block form:
$A(w_{k})=\begin{pmatrix}0_{m}&\frac{1}{2}\,(\partial_{xy}^{2}f(w_{k})-\partial_{yx}^{2}g(w_{k})^{\top})\\ \frac{1}{2}\,(\partial_{yx}^{2}g(w_{k})-\partial_{xy}^{2}f(w_{k})^{\top})&0_{n}\end{pmatrix}$. (4)
@@
As outlined in the introduction, our recent variant—Low-Rank Symplectic Gradient Adjustment (LRSGA)—proposed in Vater et al. (2025), addresses the high computational cost of explicitly forming the mixed second derivatives $\partial^{2}_{xy}f(x,y)$ and $\partial^{2}_{yx}g(x,y)$ by employing Broyden’s least–change secant updates Broyden (1965). In particular, the employed updates are the following,
$\mu_{k+1}$ $=\mu_{k}+\frac{(\partial_{x}f(w_{k+1})-\partial_{x}f(w_{k})-\mu_{k}(w_{k+1}-w_{k}))(w_{k+1}-w_{k})^{T}}{(w_{k+1}-w_{k})^{T}(w_{k+1}-w_{k})},$ (5)
$\nu_{k+1}$ $=\nu_{k}+\frac{(\partial_{y}g(w_{k+1})-\partial_{y}g(w_{k})-\nu_{k}(w_{k+1}-w_{k}))(w_{k+1}-w_{k})^{T}}{(w_{k+1}-w_{k})^{T}(w_{k+1}-w_{k})},$ (6)
obtaining $\mu_{k}\approx D(\partial_{x}f(w))=\bigl(\partial_{xx}^{2}f\,\big|\,\partial_{xy}^{2}f\bigr)\in\mathbb{R}^{m\times(m+n)}$ and $\nu_{k}\approx D(\partial_{y}g(w))=\bigl(\partial_{yx}^{2}g\,\big|\,\partial_{yy}^{2}g\bigr)\in\mathbb{R}^{n\times(m+n)}$. Henceforth we retain only the mixed–derivative block–columns, denoted $\mu_{k}]_{2}$ and $\nu_{k}]_{1}$, which approximate $\partial_{xy}^{2}f$ and $\partial_{yx}^{2}g$, respectively.
The LRSGA iterative formula is then the following:
$w_{k+1}=w_{k}-\eta\,(I-\tau\,\alpha_{k})\,F(w_{k}),$ (7)
where
$\alpha_{k}=\begin{pmatrix}0&\frac{1}{2}([\mu_{k}]_{2}-[\nu_{k}]_{1}^{\top})\\[6.0pt] \frac{1}{2}([\nu_{k}]_{1}-[\mu_{k}]_{2}^{\top})&0\end{pmatrix}\approx A(w_{k}).$ (8)
This work introduces a Limited-Memory LRSGA (LM–LRSGA) variant that preserves the theoretical framework of LRSGA while avoiding the explicit storage of the matrices $\mu_{k}$ and $\nu_{k}$. The implementation follows a two-loop recursion scheme in the spirit of L-BFGS Nocedal (1980); Liu and Nocedal (1989), leading to substantial memory savings and making the method suitable for large-scale competitive settings. Further details and evidence are provided in the subsequent sections.
3 Spectral Analysis of LRSGA
We establish a per–iteration spectral condition showing how the skew correction $(I-\tau\alpha_{k})$ stabilizes the linearized dynamics at a Nash equilibrium. The key step is an identity for the symmetric part of $(I-\tau K)H^{*}$ when $K$ is skew (Lemma 3.1); specializing $K=\alpha_{k}$ yields a spectral condition for the Jacobian of the k-frozen LRSGA map defined by $T^{\mathrm{LRSGA}}_{k}(w)=w-\eta\,(I-\tau\alpha_{k})\,F(w)$ (Proposition 3.2). Throughout, $\|\cdot\|_{2}$ denotes the spectral norm and $\mathrm{sym}(B)$ denote the symmetric part of a generic matrix B.
Lemma 3.1.
Let $H^{*}=DF(w^{*})=S^{*}+A^{*}$ with $S^{*}=S^{*\top}\succ 0$ and $A^{*}=-A^{*\top}$. For every skew matrix $K=-K^{\top}$ and $\tau>0$,
$\mathrm{sym}\!\big((I-\tau K)H^{*}\big)=S^{*}+\frac{\tau}{2}\big(S^{*}K-KS^{*}\big)\;-\;\frac{\tau}{2}\big(KA^{*}+A^{*}K\big).$ (9)
Proof.
Using $\mathrm{sym}(M)=\tfrac{1}{2}(M+M^{\top})$ and the skew-symmetry $K^{\top}=-K$ (hence $(I-\tau K)^{\top}=I+\tau K$), we have
$(I-\tau K)H^{*}=S^{*}+A^{*}-\tau KS^{*}-\tau KA^{*},\qquad H^{*\top}(I-\tau K)^{\top}=S^{*}-A^{*}+\tau S^{*}K-\tau A^{*}K,$
and summing the two identities yields (9). ∎
Proposition 3.2.
Let $w^{*}$ be a Nash equilibrium and let $H^{*}=DF(w^{*})=S^{*}+A^{*}$ with $S^{*}=S^{*\top}\succ 0$ and $A^{*}=-A^{*\top}$. Let $\alpha_{k}$ be the LRSGA skew correction at step $k$ (see (8)) and set $\Delta_{k}:=\alpha_{k}-A^{*}$ and $G_{k}:=(I-\tau\alpha_{k})H^{*}$. Consider $T^{\mathrm{LRSGA}}_{k}(w)=w-\eta\,(I-\tau\alpha_{k})\,F(w)$, whose Jacobian is $DT^{\mathrm{LRSGA}}_{k}(w^{*})=I-\eta(I-\tau\alpha_{k})H^{*}=I-\eta G_{k}$, and assume that
$0<\tau<\frac{2\,\lambda_{\min}(S^{*})}{\ \|S^{*}A^{*}-A^{*}S^{*}\|_{2}\;+\;\|S^{*}\Delta_{k}-\Delta_{k}S^{*}\|_{2}\;+\;\|\Delta_{k}A^{*}+A^{*}\Delta_{k}\|_{2}\ },$ (10)
then, there exists $\bar{\eta}_{k}>0$ such that, for every $0<\eta<\bar{\eta}_{k}$, it holds $\rho\,\!\big(D(T^{\mathrm{LRSGA}}_{k}(w^{*}))\big)\;<\;1.$
Proof.
By Lemma 3.1 with $K=\alpha_{k}=A^{*}+\Delta_{k}$,
$\mathrm{sym}(G_{k})=S^{*}+\tfrac{\tau}{2}(S^{*}A^{*}-A^{*}S^{*})-\tau A^{*2}+\tfrac{\tau}{2}\big(S^{*}\Delta_{k}-\Delta_{k}S^{*}-\Delta_{k}A^{*}-A^{*}\Delta_{k}\big).$
Since $\tau(A^{*})^{2}\succeq 0$ we obtain
$\mathrm{sym}(G_{k})\ \succeq\ c_{k}I,\quad\text{with}\quad c_{k}:=\lambda_{\min}(S^{*})-\frac{\tau}{2}\Big(\|S^{*}A^{*}-A^{*}S^{*}\|_{2}+\|S^{*}\Delta_{k}-\Delta_{k}S^{*}\|_{2}+\|\Delta_{k}A^{*}+A^{*}\Delta_{k}\|_{2}\Big).$
The condition (10) implies $c_{k}>0$. If $\lambda$ is an eigenvalue of $G_{k}$, then
$|1-\eta\lambda|^{2}\;\leq\;1-2\eta\,\Re\lambda+\eta^{2}|\lambda|^{2}\;\leq\;1-2\eta c_{k}+\eta^{2}\|G_{k}\|_{2}^{2}.$
Setting
$\bar{\eta}_{k}:=\frac{2c_{k}}{\|G_{k}\|_{2}^{2}}>0,$
we obtain $|1-\eta\lambda|<1$ for every $0<\eta<\bar{\eta}_{k}$, hence $\rho(I-\eta G_{k})<1$ and therefore $\rho(DT^{\mathrm{LRSGA}}_{k}(w^{*}))<1$. ∎
As proved in Vater et al. (2025), letting $w^{*}\in\Omega$ be a NE point and assuming there is a $\delta>0$ such that $\|{\mu_{k}-D(\partial_{x}f(w^{*}))}\|_{2}\leq\delta$ and $\|{\nu_{k}-D(\partial_{y}g(w^{*}))}\|_{2}\leq\delta$ for $\mu_{k}$ and $\nu_{k}$ (see equations (5) and (6)), it holds that the matrix $\alpha_{k}$, defined in (8), satisfies the inequality $\|{\Delta_{k}}\|_{2}\leq\delta$ with $\Delta_{k}:=\alpha_{k}-A^{*}$. Exploiting this result, the $\tau$-bound per iteration in Proposition 3.2 can be reformulated in global terms as follows.
Corollary 3.3.
Let $w^{*}\in\Omega$ and assume there is a $\delta>0$ such that it holds $\|{\mu_{k}-D(\partial_{x}f(w^{*}))}\|_{2}\leq\delta$ and $\|{\nu_{k}-D(\partial_{y}g(w^{*}))}\|_{2}\leq\delta$. Considering again $T^{\mathrm{LRSGA}}_{k}(w)=w-\eta\,(I-\tau\alpha_{k})\,F(w)$, if
$0<\tau<\frac{2\lambda_{\min}(S^{*})}{\|S^{*}A^{*}-A^{*}S^{*}\|_{2}+2(\|S^{*}\|_{2}+\|A^{*}\|_{2})\,\delta},$
then there exists $\bar{\eta}_{k}>0$ such that, for every $0<\eta<\bar{\eta}_{k}$, it holds $\rho\,\big(DT^{\mathrm{LRSGA}}_{k}(w^{*})\big)\;<\;1.$
Proof.
Simply by calculus we have
$\|S^{*}\Delta_{k}-\Delta_{k}S^{*}\|_{2}\leq 2\|S^{*}\|_{2}\|\Delta_{k}\|_{2},\qquad\|\Delta_{k}A^{*}+A^{*}\Delta_{k}\|_{2}\leq 2\|A^{*}\|_{2}\|\Delta_{k}\|_{2},\qquad$
By hypothesis we can apply Vater et al. (2025) result about $\Delta_{k}$ obtaining:
$\|S^{*}\Delta_{k}-\Delta_{k}S^{*}\|_{2}\leq 2\|S^{*}\|_{2}\delta,\qquad\|\Delta_{k}A^{*}+A^{*}\Delta_{k}\|_{2}\leq 2\|A^{*}\|_{2}\delta,$
and then the $\tau$–condition in Proposition 3.2 is implied by the global condition
$0<\tau<\frac{2\lambda_{\min}(S^{*})}{\|S^{*}A^{*}-A^{*}S^{*}\|_{2}+2(\|S^{*}\|_{2}+\|A^{*}\|_{2})\,\delta}$
We remark that these results specialize to SGA in the limit $\Delta_{k}\!\to\!0$. In particular, setting $\Delta_{k}=0$ yields
$\rho\,\!\bigl(DT^{\mathrm{SGA}}_{\star}(w^{*})\bigr)<1\quad\text{for}\quad T^{\mathrm{SGA}}_{\star}(w):=w-\eta\,(I-\tau A^{*})\,F(w),$
under the same restriction on $\tau$ and a corresponding choice of $\eta$.
These stability estimates, stated for the full LRSGA skew correction matrix $\alpha_{k}$, will be used in Section 5 to prove local convergence of LM–LRSGA, since the limited-memory skew correction matrix falls within the perturbation framework of Proposition 3.2 (via Corollary 3.3).
4 Limited Memory LRSGA
We begin this section by introducing a further ’simplification’ for LRSGA before address the problem the storage difficulties. So, instead of working directly on LRSGA, we will work on this revised version.
@@ 4 Limited Memory LRSGA
It is algebraically and computationally equivalent (at the level of per–iteration arithmetic) to (i) form $\mu_{k}$ and $\nu_{k}$ and then extract only the mixed blocks $\mu_{k}]_{2}$ and $\nu_{k}]_{1}$, or (ii) update directly the mixed blocks. Indeed, starting from (5 and 6), by partitioning $\mu_{k}=\bigl([\mu_{k}]_{1}\,\big|\,[\mu_{k}]_{2}\bigr)$ and $\nu_{k}=\bigl([\nu_{k}]_{1}\,\big|\,[\nu_{k}]_{2}\bigr)$, one readily verifies
@@
where $f_{k}:=f(w_{k})$ and $g_{k}:=g(w_{k})$. Thus, in either formulation the same quantities—namely the products $\mu_{k}]_{1}(x_{k+1}-x_{k})$ and $\nu_{k}]_{2}(y_{k+1}-y_{k})$—are required, and the per–step computational cost is comparable.
In a neighborhood of a stable Nash equilibrium $w^{*}$, the second–order necessary conditions imply that the own–player Hessian blocks are positive semidefinite,
@@
Locally around $w^{*}$, we assume that the diagonal blocks are dominated by their coercive (positive) component, eventually corresponding to a stabilizing quadratic behavior of the form
$f(x,y)+\tfrac{\lambda_{x}}{2}\|x-x^{*}\|^{2},\qquad g(x,y)+\tfrac{\lambda_{y}}{2}\|y-y^{*}\|^{2},$
with $\lambda_{x},\lambda_{y}>0$. Such contributions do not alter rotational effects in the local dynamics.
Since our objective is to eliminate the rotational component of the game dynamics, we deliberately neglect the diagonal Hessian blocks and retain only the off–diagonal coupling, which is responsible for non-potential behavior. Accordingly, we approximate the local Jacobian $H(w_{k})$ by its skew–symmetric part, while assuming that the diagonal contribution is implicitly regularized by the underlying positive curvature.
To this end, we introduce constant parameters $\varepsilon_{x},\varepsilon_{y}\geq 0$, interpreted as dominant positive eigenvalues of the diagonal blocks, and rewrite the secant conditions for the off–diagonal approximations as
$M_{k+1}$ $=M_{k}+\frac{\big(\partial_{x}f_{k+1}-\partial_{x}f_{k}-\varepsilon_{x}(x_{k+1}-x_{k})-M_{k}(y_{k+1}-y_{k})\big)(y_{k+1}-y_{k})^{\top}}{(w_{k+1}-w_{k})^{\top}(w_{k+1}-w_{k})},$ (13)
$N_{k+1}$ $=N_{k}+\frac{\big(\partial_{y}g_{k+1}-\partial_{y}g_{k}-\varepsilon_{y}(y_{k+1}-y_{k})-N_{k}(x_{k+1}-x_{k})\big)(x_{k+1}-x_{k})^{\top}}{(w_{k+1}-w_{k})^{\top}(w_{k+1}-w_{k})}.$
The terms $\varepsilon_{x}(x_{k+1}-x_{k})$ and $\varepsilon_{y}(y_{k+1}-y_{k})$ model the dominant coercive effect of the diagonal blocks and act as a built–in Levenberg–Marquardt–type regularization, not affecting the rotational structure of the dynamics.
Remark 4.1.
The introduction of $\varepsilon_{x},\varepsilon_{y}$ in (13) should be understood as a surrogate for the diagonal (own–player) blocks that are implicitly present in the original secant relations. Indeed, when one derives the mixed-block recursions induced by the full least–change updates, the formulas contain the diagonal-block actions $\mu_{k}]_{1}(x_{k+1}-x_{k})$ and $\nu_{k}]_{2}(y_{k+1}-y_{k})$. The replacements
$\mu_{k}]_{1}(x_{k+1}-x_{k})\approx\varepsilon_{x}(x_{k+1}-x_{k}),\qquad[\nu_{k}]_{2}(y_{k+1}-y_{k})\approx\varepsilon_{y}(y_{k+1}-y_{k}),$
provide the minimal block-level closure that preserves a secant-type update for the mixed blocks while avoiding the explicit formation and storage of the full matrices $\mu_{k}\in\mathbb{R}^{m\times(m+n)}$ and $\nu_{k}\in\mathbb{R}^{n\times(m+n)}$. In this way one can update and store directly $M_{k}\in\mathbb{R}^{m\times n}$ and $N_{k}\in\mathbb{R}^{n\times m}$ (instead of computing $\mu_{k},\nu_{k}$ and then extracting only $\mu_{k}]_{2}$ and $\nu_{k}]_{1}$), while still accounting for the dominant coercive effect of the diagonal blocks through $\varepsilon_{x},\varepsilon_{y}$. Finally, this approximation does not change the equilibrium conditions of the iteration: since $\alpha_{k}$ is skew-symmetric, $I-\tau\alpha_{k}$ is nonsingular for every $\tau>0$, hence $(I-\tau\alpha_{k})F(w_{k})=0\iff F(w_{k})=0$.
As a result, the correction matrix adopted in this slightly revised LRSGA is purely skew–symmetric,
$\overline{\alpha}_{k}=\begin{pmatrix}0&\tfrac{1}{2}(M_{k}-N_{k}^{\top})\\[6.0pt] \tfrac{1}{2}(N_{k}-M_{k}^{\top})&0\end{pmatrix}$, (14)
which isolates and counteracts the rotational component of the game Jacobian, while the stabilizing diagonal curvature is implicitly absorbed into the step regularization.
@@
In any case, as highlighted, LRSGA requires storing the matrices $M_{k}$ and $N_{k}$ to approximate the mixed second derivatives, which can be memory–intensive for high–dimensional models. To address this, we propose a Limited-Memory LRSGA variant that preserves LRSGA’s theoretical robustness while substantially reducing memory usage. Our strategy follows the two-loop recursion method, successfully employed by L-BFGS Nocedal (1980); Liu and Nocedal (1989). The key idea is to avoid storing the full updated matrices $M_{k}$ and $N_{k}$; instead, we retain only the most recent $\ell$ pairs among the four vectors $\{s_{k}^{x},s_{k}^{y},y_{k}^{f},y_{k}^{g}\}$, defined as
$s_{k}^{x}=x_{k+1}-x_{k},\qquad s_{k}^{y}=y_{k+1}-y_{k},$
$y_{k}^{f}=\partial_{x}f_{k+1}-\partial_{x}f_{k}-\varepsilon_{x}\,s_{k}^{x},\qquad y_{k}^{g}=\partial_{y}g_{k+1}-\partial_{y}g_{k}-\varepsilon_{y}\,s_{k}^{y}.$